# An Efficient Closed-Form Solution to Full Visual-Inertial State Initialization

### v1

@@ V-A Closed-form initialization from linear point constraints
Once $\mathtt{R}_{ij}$ and $\mathbf{b}^{g}$ are initialized, the relation between inertial measurements and image tracks can be expressed linearly. Let the sets $\mathcal{Z}_{i}=\{\mathbf{z}_{i}^{1},\ldots,\mathbf{z}_{i}^{n},\ldots,\mathbf{z}_{i}^{N}\}$ and $\mathcal{Z}_{j}=\{\mathbf{z}_{j}^{1},\ldots,\mathbf{z}_{j}^{n},\ldots,\mathbf{z}_{j}^{N}\}$ denote the image coordinates of NN visual features extracted in an image at $t_{i}$ and tracked up to a later image, not necessarily consecutive, taken at $t_{j}$. Assuming a pinhole camera model, every feature $\mathbf{z}_{c}^{n}\in\Omega\subset\mathbb{R}^{2}$ in both images $c\in\{i,j\}$ can be backprojected to a 3D point $\mathbf{p}_{c}^{n}\in\mathbb{R}^{3}$ in the local camera frame as follows
@@
The rotation $\mathtt{R}_{i\tau}\in{SO}(3)$ comes from the IMU preintegration from time $t_{i}$ up to time $\tau$ using Eq. 1. Note that $\mathbf{b}^{g}$ was already initialized from the first two frames. The vector $\tilde{\mathbf{a}}_{\tau}$ contains the accelerometer readings. By stacking the constraints from the NN tracked features from frame $i$ to frame $j$, we obtain the overdetermined linear system $\boldsymbol{\Xi}\,\mathbf{x}=\mathbf{s}$, which yields the following least-squares problem $\arg\min_{\mathbf{x}}\left(\boldsymbol{\Xi}\,\mathbf{x}-\mathbf{s}\right)^{2}\rightarrow\hat{\mathbf{x}}=\mathbf{v}_{\text{min}}\left(\boldsymbol{\Xi}^{\top}\boldsymbol{\Xi}\right)$ (20) which we solve via singular value decomposition (SVD), being the optimal solution $\hat{\mathbf{x}}$ the eigenvector associated to the smallest eigenvalue $\mathbf{v}_{\text{min}}\left(\boldsymbol{\Xi}^{\top}\boldsymbol{\Xi}\right)$. Each row of the matrix $\boldsymbol{\Xi}$ corresponds to one instance of Eq. (18), leading to a total of 3(M1)N3(M-1)N equations and $9+NM$ unknowns, being $M$ the number of images considered. $\boldsymbol{\Xi}$ denotes the coefficients matrix

### v2

@@ V-A Closed-form initialization from linear point constraints
Once $\mathtt{R}_{ij}$ and $\mathbf{b}^{g}$ are initialized, the relation between inertial measurements and image tracks can be expressed linearly. Let the sets $\mathcal{Z}_{i}=\{\mathbf{z}_{i}^{1},\ldots,\mathbf{z}_{i}^{n},\ldots,\mathbf{z}_{i}^{N}\}$ and $\mathcal{Z}_{j}=\{\mathbf{z}_{j}^{1},\ldots,\mathbf{z}_{j}^{n},\ldots,\mathbf{z}_{j}^{N}\}$ denote the image coordinates of $N$ visual features extracted in an image at $t_{i}$ and tracked up to a later image, not necessarily consecutive, taken at $t_{j}$. Assuming a pinhole camera model, every feature $\mathbf{z}_{c}^{n}\in\Omega\subset\mathbb{R}^{2}$ in both images $c\in\{i,j\}$ can be backprojected to a 3D point $\mathbf{p}_{c}^{n}\in\mathbb{R}^{3}$ in the local camera frame as follows
@@
The rotation $\mathtt{R}_{i\tau}\in{SO}(3)$ comes from the IMU preintegration from time $t_{i}$ up to time $\tau$ using Eq. 1. Note that $\mathbf{b}^{g}$ was already initialized from the first two frames. The vector $\tilde{\mathbf{a}}_{\tau}$ contains the accelerometer readings. By stacking the constraints from the $N$ tracked features from frame $i$ to frame $j$, we obtain the overdetermined linear system $\boldsymbol{\Xi}\,\mathbf{x}=\mathbf{s}$, which yields the following least-squares problem $\arg\min_{\mathbf{x}}\left(\boldsymbol{\Xi}\,\mathbf{x}-\mathbf{s}\right)^{2}\rightarrow\hat{\mathbf{x}}=\mathbf{v}_{\text{min}}\left(\boldsymbol{\Xi}^{\top}\boldsymbol{\Xi}\right)$ (20) which we solve via singular value decomposition (SVD), being the optimal solution $\hat{\mathbf{x}}$ the eigenvector associated to the smallest eigenvalue $\mathbf{v}_{\text{min}}\left(\boldsymbol{\Xi}^{\top}\boldsymbol{\Xi}\right)$. Each row of the matrix $\boldsymbol{\Xi}$ corresponds to one instance of Eq. (18), leading to a total of 3(M1)N3(M-1)N equations and $9+NM$ unknowns, being $M$ the number of images considered. $\boldsymbol{\Xi}$ denotes the coefficients matrix